ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 238, 2020

Balázs. B. A. and Sz. Mészáros The minimal degree standard identity on M_{n}E^{2} and M_{n}E^{3}

279

Barriga, E. Definably compact groups definable in real closed fields  121

Berasategui, M. and D. Carando Unconditional Schauder frames of translates in L_{p}(ℝ^{d})

687

Bès, J. and D. Papathanasiou Algebrable sets of hypercyclic vectors for convolution operators

91

Bucić, M., M. Kwan, A. Pokrovskiy, B. Sudakov, T. Tran and A. Z. Wagner Nearly-linear monotone paths in edge-ordered graphs

663

Buijs, U., Y. Félix, A. Murillo and D. Tanré Lie models of simplicial sets and representability of the Quillen functor

313

Bukh, B. and C. Cox Nearly orthogonal vectors and small antipodal spherical codes

359

Bürger, C. and M. Pitz Partitioning edge-coloured infinite complete bipartite graphs into monochromatic paths

479

Callegaro, F. and M. Salvetti Families of superelliptic curves, complex braid groups and generalized Dehn twists

945

Carando, D. See M. Berasategui Chudnovsky, M., A. Scott, P. Seymour and S. Spirkl Proof of

the Kalai–Meshulam conjecture

639

Cox, C. See B. Bukh Damjanović, D. and D. Xu On classification of higher rank Anosov

actions on compact manifold

745

Engel, K., T. Mitsis, C. Pelekis and C. Reiher Projection inequalities for antichains

61

Félix, Y. See U. Buijs Ghara, S. The orbit of a bounded operator under the Möbius group

modulo similarity equivalence

167

Gogolev, A., B. Kalinin and V. Sadovskaya Local rigidity of Lyapunov spectrum for toral automorphisms

389

Halawi, H. and A. Segal The degenerate principal series representations of exceptional groups of type E_{6} over p-adic fields

537

He, W. Random walks on linear groups satisfying a Schubert condition  593

Huang, L., J. Wu and J. Xu Metric properties of the product of consecutive partial quotients in continued fractions

901

Jiao, Y., F. Sukochev and D. Zanin On subspaces spanned by freely independent random variables in noncommutative L_{p}-spaces

431

Jung, J. Bounding the number of nodal domains of eigenfunctions without singular points on the square

1

Kalinin, B. See A. Gogolev Kwan, M. See M. Bucić Lechner, R., P. Motakis, P. F. X. Müller and

T. Schlumprecht Strategically reproducible bases and the factorization property

13

Lieberman, M., J. Rosický and S. Vasey Sizes and filtrations in accessible categories

243

Margolis, L. and O. Schnabel The twisted group ring isomorphism problem over fields

209

Mészáros, Sz. See B. A. Balázs Mitsis, T. See K. Engel Moshchevitin, N., B. Murphy and I. Shkredov Popular products

and continued fractions

807

Motakis, P. See R. Lechner Müller, P. F. X. See R. Lechner Murillo, A. See U. Buijs Murphy, B. See N. Moshchevitin Nie, H. and K. M. Pilgrim Boundedness of hyperbolic components

of Newton maps

837

Oh, J. S. and Q. Zhong On Erdős–Ginzburg–Ziv inverse theorems for dihedral and dicyclic groups

715

Ombrosi, S., C. Pérez, E. Rela and I. P. R. os A note on generalized Fujii–Wilson conditions and BMO spaces

571

Pérez, C. See S. Ombrosi Papathanasiou, D. See J. Bès Pelekis, C. See K. Engel Pilgrim, K. M. See H. Nie Pitz, M. See C. Bürger Pokrovskiy, A. See M. Bucić Reiher, C. See K. Engel Rela, E. See S. Ombrosi Rosický, J. See M. Lieberman Sadovskaya, V. See A. Gogolev Salvetti, M. See F. Callegaro Schlumprecht, T. See R. Lechner Schnabel, O. See L. Margolis Scott, A. See M. Chudnovsky Segal, A. See H. Halawi Seymour, P. See M. Chudnovsky Shkredov, I. See N. Moshchevitin Spirkl, S. See M. Chudnovsky Sudakov, B. See M. Bucić Sukochev, F. See Y. Jiao

Suzuki, M. Quaternion distinguished generic representations of GL_{2n}  871

Tanré, D. See U. Buijs Tărnăuceanu, M. Detecting structural properties of finite groups

by the sum of element orders

629

Tran, T. See M. Bucić Vasey, S. See M. Lieberman Wagner, A. Z. See M. Bucić Wang, J. and X. Yao On Julia limiting directions of meromorphic

functions

405

Woods, W. On the catenarity of virtually nilpotent mod-p Iwasawa algebras

501

Wu, J. See L. Huang Xu, D. See D. Damjanović Xu, J. See L. Huang Yao, X. See J. Wang Zanin, D. See Y. Jiao Zhong, Q. See J. S. Oh os, I. P. R. See S. Ombrosi